OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is 3x – 5 and whose width is x-4

OpenStudy (whpalmer4):

The area of a rectangle is width * length. What do you get when you multiply those two quantities?

OpenStudy (anonymous):

20

OpenStudy (whpalmer4):

Unless you have additional information (the value of x) that you didn't provide, that isn't the answer. length = 3x-5 width = x-4 length * width = (3x-5)(x-4)

OpenStudy (anonymous):

3x^2-17x+20